A man goes 150m due east and then 200m due north.How far is he starting from the point?

Grade:10

1 Answers

Ishaan Jain
6 years ago
Let east direction be î and north direction be j cap The final position vector therefore is150 î + 200 jAnd magnitude is √ (150)^2 + (200)^2That is 250 m Direction of man from initial position isTan # = 150/200 = 3/4 = 37°The man is 250 m and 37 ° North of east from his initial position